OpenStudy (anonymous):

partial fraction decomposition: (2x^3+4x-1)/(x^2+2)^2

OpenStudy (anonymous):

\[\frac{2x^3+4x-1}{(x^2+2)^2}=\frac{Ax+B}{x^2+2}+\frac{Cx+D}{(x^2+2)^2}\] or you can go the longer way using \[\frac{2x^3+4x-1}{(x^2+2)^2}=\frac{Ax+B}{x^2+2}+\frac{Cx^3+Dx^2+Ex+F}{(x^2+2)^2}\] I'll refer to the first way. \[\begin{align*}\frac{2x^3+4x-1}{(x^2+2)^2}&=\frac{Ax+B}{x^2+2}+\frac{Cx+D}{(x^2+2)^2}\\\\ 2x^3+4x-1&=(Ax+B)(x^2+2)+(Cx+D)\\\\ 2x^3+4x-1&=Ax^3+Bx^2+2Ax+2B+Cx+D\\\\ 2x^3+4x-1&=Ax^3+Bx^2+(2A+C)x+(2B+D) \end{align*}\] yielding the system \[\begin{cases} A=2\\B=0\\2A+C=4\\2B+D=-1 \end{cases}~~\Rightarrow~~A=2,~B=0,~C=0,~D=-1\]